# source.bib

### v1

@@
Skelet #17 and the determination of BB(5)
@@ 1.1. Introduction
A massive undertaking by The Busy Beaver Challenge team [bbchallenge] has narrowed the determination of $\text{BB}(5)$ down to showing non-halting of a handful of machines. While most of the machines can be grouped into categories, the machine dubbed “Skelet #17” [skelet17] has eluded all known attempts at classification. The goal of this paper is to prove that Skelet #17 does not halt (Theorem 1.3).
@@ 1.3. Formalization
The work of this paper has recently been formalized in Coq [formalproof], along with everything else in the $\text{BB}(5)$ program. As a result, we are able to conclude:
@@ 3.2. Empty and embanked states
    $a_{1}(E)<3\cdot 2^{2k}-1-\overline{a_{0}(E)}$,
where as in Definition 2.1, we denote $\overline{a_{0}(E)}\in\{0,1\}$ as the mod $2$ reduction of $a_{0}(E)$.
@@
The parity of $a_{0}(E)$ affects the simplification of the above inequality: namely, it is equivalent to
$\begin{cases}a_{1}(E)+\frac{a_{0}(E)+1}{2}+1<2^{2k+2}-1-\left(2^{2k}-\frac{a_{0}(E)+1}{2}\right)&a_{0}(E)\equiv 1\bmod 2\\ a_{1}(E)+\frac{a_{0}(E)}{2}+1<2^{2k+2}-1-\left(2^{2k}-\frac{a_{0}(E)+2}{2}\right)&a_{0}(E)\equiv 0\bmod 2,\end{cases}$
which simplifies to
$a_{1}(E)<\begin{cases}3\cdot 2^{2k}-2&a_{0}(E)\equiv 1\bmod 2\\ 3\cdot 2^{2k}-1&a_{0}(E)\equiv 0\bmod 2.\end{cases}$
This is exactly condition (2), so we win. ∎
@@ 4.2. Endgame
Our setup for the end of the proof is
$S_{k}\to E\to E^{\prime}\to E^{\prime\prime},$
where $E$, $E^{\prime}$ and $E^{\prime\prime}$ are all rooted embanked states satisfying $h(N^{-1}(E))=(2^{2k}-3,2^{2k}+2)$, $h(N^{-1}(E^{\prime}))=(1,2^{2k+1}-2)$ and $h(N^{-1}(E^{\prime\prime}))=(1,2^{2k+1}-1)$. In particular, we have $E^{\prime}\to E^{\prime\prime}$ since Proposition 4.1(2) guarantees that $E^{\prime}$ will satisfy the bounds in Proposition 3.4 so that $E^{\prime}$ is weakly embanked, and then Corollary 3.8 implies that $E^{\prime}$ is embanked. Also, we have $E=(N^{\prime})^{5}(S_{k})$. An explicit computation yields

### v2

@@
Skelet #17 and the fifth Busy Beaver number
@@ 1.1. Introduction
A massive undertaking by The Busy Beaver Challenge team [bbchallenge] has narrowed down the determination of $\text{BB}(5)$ to showing non-halting of approximately 30 “holdout” machines [bbchallengeholdouts]. Of those machines, there is significant overlap with a different list of 43 holdouts for which Georgi “Skelet” Georgiev claimed to find in 2003 [skelethomepage]: the machine dubbed Skelet #17 [skelet17] is the focus of our paper.
Although most of the 30 holdouts can be grouped into well-known categories, analysis of Skelet #17 has eluded all known attempts at classification. Members of [bbchallenge] are in general agreement that Skelet #17, along with Skelet #1, comprise the two most difficult holdouts to analyze. In this paper, we show that Skelet #17 does not halt (Theorem 1.3).
@@ 1.3. Formalization
The author originally wrote up a sketch of this paper in 2023. Recently, “Mxdys” has adapted its contents to a formal proof of Skelet #17’s nonhalting [formalproof17], along with formalizing everything else in the $\text{BB}(5)$ program [formalproof]. As a result, we are able to conclude:
@@ 3.2. Empty and embanked states
    $a_{1}(E)<3\cdot 2^{2k}-1$.
@@
Note that $a_{0}(E)$ is necessarily even (since $E$ is empty), so the above inequality becomes
$a_{1}(E)+\frac{a_{0}(E)}{2}+1$ $<2^{2k+2}-1-\left(2^{2k}-\frac{a_{0}(E)+2}{2}\right),$
which further simplifies to $a_{1}(E)<3\cdot 2^{2k}-1.$ This is exactly condition (2), so we win. ∎
@@ 4.2. Endgame
Our setup for the end of the proof is the sequence of state transitions $S_{k}\to E\to E^{\prime}\to E^{\prime\prime}$, where $E$, $E^{\prime}$ and $E^{\prime\prime}$ are all rooted embanked states satisfying $h(N^{-1}(E))=(2^{2k}-3,2^{2k}+2)$, $h(N^{-1}(E^{\prime}))=(1,2^{2k+1}-2)$ and $h(N^{-1}(E^{\prime\prime}))=(1,2^{2k+1}-1)$. In particular, we have $E^{\prime}\to E^{\prime\prime}$ since Proposition 4.1(2) guarantees that $E^{\prime}$ will satisfy the bounds in Proposition 3.4 so that $E^{\prime}$ is weakly embanked, and then Corollary 3.8 implies that $E^{\prime}$ is embanked. Also, we have $E=(N^{\prime})^{5}(S_{k})$. An explicit computation yields